### v1

# CLEAR-IR: Clarity-Enhanced Active Reconstruction of Infrared Imagery

Abstract
This paper presents a novel approach for enabling robust robotic perception in dark environments using infrared (IR) stream. IR stream is less susceptible to noise than RGB in low-light conditions. However, it is dominated by active emitter patterns that hinder high-level tasks such as object detection, tracking and localisation. To address this, a U-Net-based architecture is proposed that reconstructs clean IR images from emitter-populated input, improving both image quality and downstream robotic performance. This approach outperforms existing enhancement techniques and enables reliable operation of vision-driven robotic systems across illumination conditions from well-lit to extreme low-light scenes.
I Introduction
Lighting-invariant vision systems are desirable for enabling robots to operate robustly across diverse and unpredictable environments without requiring modifications to the underlying perception pipeline. In order to support high-level tasks such as object detection, semantic segmentation, and image classification, the vision system must remain reliable even in low light or completely dark scenes. Such capabilities are critical in domains like mine shaft exploration, post-disaster victim identification, nuclear facility inspection, and visual loop closure in feature-deprived environments using aruco markers.
Conventional RGB cameras, while effective under well-lit conditions, suffer from significant degradation in low-light scenes [1]. Image quality is compromised due to sensor-level limitations, such as dark current and readout noise. Attempts to compensate using an increased exposure time introduce motion blur, rendering them ineffective for dynamic environments. Furthermore, data-driven approaches that require training networks in long-exposure or low-light image pairs add complexity and reduce generalisability across different camera models. In addition, such methods are fundamentally limited by information-theoretic constraints and if critical scene information is not captured by the sensor due to noise or extreme darkness, no amount of training data can fully reconstruct it. The effectiveness of these networks is therefore heavily dependent on both data availability and inherent scene observability.
Another approach involves equipping robots with onboard illumination systems [2]. However, in certain rich environments such as underground tunnels or disaster zones, the Tyndall effect causes the light to scatter off airborne particles, saturating the camera view with undesirable visual artefacts and occluding key features in the scene. This method also suffers from overexposure, specular reflection, and increased power consumption. Moreover, the use of active illumination can produce strong shadows that obscure structural details and complicate downstream vision tasks such as object detection and localisation. To address these challenges, this work explores the use of infrared (IR) imaging, particularly active IR systems that operate independently of ambient illumination. As seen in Figure 1, the dark scene captured by the RGB camera fails to present good visual cues, whereas the same scene in the IR spectrum can be visualised with several objects in the scene. Although IR cameras are promising for dark-scene perception, they are not without their own limitations. Passive IR sensors, which operate without emitters, avoid structured-light artefacts but are typically expensive and limited to low-resolution outputs, making them impractical for robotics. Active IR systems, introduce structured light patterns that degrade image quality, interfere with feature extraction, and disrupt high-level perception tasks such as object detection, tracking, and SLAM.
To overcome this, a deep learning-based framework, CLEAR-IR (Clarity-Enhanced Active Reconstruction of Infrared Imagery), is proposed. The framework is designed to reconstruct clean and interpretable IR images from raw emitter-populated inputs. The method leverages a U-Net architecture trained to suppress structured light noise while preserving spatial and textural details critical for vision-based tasks. The output is a denoised, visually coherent IR image that can serve as a robust input to standard perception modules. This is the first work to leverage the infrared stream for high-level robotic tasks by removing active emitter patterns in scenes ranging from well-illuminated to extreme low-light. The contributions of this work are:
  • CLEAR-IR, a U-Net–based framework tailored for reconstructing clean IR images by suppressing structured-light artefacts, is introduced.
  • A composite loss function is developed to preserve spatial and textural details critical for robotic perception while robustly removing high-frequency emitter patterns.
  • Evaluation of IR image reconstruction for robotics is presented, demonstrating significant improvements across key tasks such as feature matching, object detection, and marker detection.
(a) Low-light room captured using an RGB camera
(b) Low-light room captured using an IR camera
Figure 1: Low-light room captured using different cameras
II Related Work
Vision-based robotic perception in challenging lighting conditions has been an active area of research. One line of work has focused on improving the performance of standard RGB cameras in low light. This includes sensor-level enhancements such as improved read-out noise reduction and pixel architectures [3], as well as computational photography techniques such as multiexposure fusion [4] and burst photography [5]. Although these techniques can improve image quality, they are often computationally intensive and can introduce motion blur or require static scenes, limiting their use in dynamic robotic applications.
A separate body of work leverages deep learning for low-light image enhancement. Prominent examples include supervised learning methods that train neural networks on paired low light and well lit images [6, 7]. These approaches, however, are highly dependent on the quality and diversity of the training data and may not generalise well to different camera hardware or unseen environments. Unsupervised and semi-supervised methods have also been explored to mitigate the need for paired data as seen in [8, 9, 10], but they often struggle to remove complex artefacts such as structured light patterns found in active IR systems, which this work explores. Generative adversarial networks (GANs) have been used for enhancement in low light as inferred from [11, 12, 13], but these methods do not translate well to extremely low light scenes, which require some form of active sensing. Moreover, the artefacts introduced by these methods are particularly problematic for robotics: vision-based SLAM pipelines may incorrectly latch onto these artificial structures as features, leading to spurious correspondences, tracking drift, and eventual map corruption. Thus, preserving fine-grained but genuine details is critical for ensuring robustness in downstream tasks such as vSLAM and visual odometry.
Active infrared (IR) systems, which operate by illuminating a scene with an IR light source and capturing the reflected light, have been a viable alternative for dark-scene perception. These systems have been widely adopted in applications such as facial recognition [14] and depth sensing [15]. However, the structured light patterns they use for scene illumination and depth estimation, while beneficial for some tasks, present a significant challenge for 2D vision tasks like object detection and segmentation. Repetitive high-frequency patterns can be misinterpreted as texture or objects by standard vision algorithms, leading to false positives or missed detections, which limits the amount of work done using these sensors. Previous works have only attempted to remove fixed patterns from IR images, as seen in [16, 17, 18]. This just enhances the extremely noisy input of IR images, but does not focus on making it more useful for high-level robotics tasks.
The U-Net architecture, with its symmetric encoder-decoder structure and skip connections, has shown remarkable success in a variety of image-to-image translation tasks, and was first introduced for biomedical image segmentation in [19]. Its ability to capture both high-level semantic information from the encoder path and fine-grained spatial details from the decoder path makes it an ideal candidate for the specific denoising problem encountered in IR images. This architecture has also been successfully adapted for other image reconstruction tasks, such as super-resolution [20] and general image denoising [21].
In contrast to existing methods that primarily focus on low-light enhancement or general noise removal, this work addresses the specific and unique problem of structured light artefacts in active IR imagery. Using the U-Net architecture, it is possible to effectively learn to suppress these specific patterns while preserving the crucial underlying semantic information for high-level vision tasks.
III Methodology
This section presents the CLEAR-IR framework. This method leverages a U-Net architecture and a composite loss function to effectively remove structured light artefacts while preserving critical visual details.
III-A Emitter Pattern Characteristics
Active infrared (IR) systems enable vision in dark environments by projecting an illumination pattern onto the scene. This pattern is often a pseudo-random pattern which is a semi-deterministic and repeatable sequence of dots generated by an algorithm. The raw IR image captured by the camera is, therefore, populated with these dots, which can be misidentified as texture or objects by standard vision algorithms, complicating tasks like object detection and segmentation.
The intensity of these dots in the captured image is not uniform and depends on four main factors:
  • Object Reflectivity: Objects with higher reflectivity in the IR spectrum will appear brighter, as they reflect a sharper emitter pattern to the camera, as shown in Figure2a, where the plastic screen reflects more emitter pattern than the black foam.
  • Distance to the Emitter: The intensity of the light from the emitter decreases with the square of the distance, following the inverse square law. Consequently, closer objects will show a more intense and prominent dot pattern, where the patterns are larger when closer and smaller when they move further away from the camera, as seen in Figure2b and Figure2a, where the closer part of the sheet has more shaper patterns than the further part.
  • Emitter Power: The overall intensity of the pattern is directly proportional to the laser power of the emitter, where a higher power leads to a denser, more defined pattern. This is also a critical factor in the appearance of the pattern, as seen in Figure 2c, where a lower laser power prevents the room from being seen while a larger laser power populates the room with the emitter pattern
  • Room Brightness: This factor contributes to the emitter power configuration, where the same emitter pattern may appear dimmer in well-illuminated scenes while appearing sharper in low-light scenes, as seen in Figure2d and Figure2e, where ambient light affects the emitter visibility by dimming the laser power.
(a) Emitter pattern’s appearance based on object reflectivity
(b) Emitter pattern’s appearance based on distance
(c) Changes in the emitter pattern’s intensity based on the laser power of the emitter
(d) Emitter pattern in a bright room
(e) Emitter pattern in a dark room
Figure 2: The key factors influencing the appearance of the structured light emitter pattern in active infrared imagery.
This variability in pattern appearance across different scenes creates a significant challenge in selecting adequate losses that can remove them from the images while being scene-agnostic.
III-B U-Net Architecture for Emitter Pattern Removal
To effectively reconstruct the image and remove the noise caused by the emitter pattern, a standard U-Net architecture was used. This convolutional neural network is particularly well suited for image-to-image translation tasks due to its symmetric encoder-decoder structure and strategic use of skip connections, as depicted in Figure 3.
The network is configured to process greyscale images with an input resolution of 640x480 pixels. The encoder path progressively downsamples the image, using a series of convolutional and max-pooling layers to extract increasingly abstract features and high-level contextual information.
The decoder path then reconstructs the image using transposed convolutional layers to upsample the features back to the original resolution. A crucial element of this architecture are its skip connections, which concatenate the feature maps from the encoder to the corresponding layers in the decoder. These connections are vital for preserving fine-grained spatial details that might otherwise be lost during the downsampling process. The network combines these features at each upsampling stage to produce a final, single-channel output image that is the same size as the input, with the emitter pattern effectively removed.
Figure 3: The U-Net architecture used for image reconstruction and emitter pattern removal.
III-C Dataset and Data Augmentation
A custom paired dataset of IR images and their corresponding clean ground-truth RGB images converted to greyscale was created for model training using an Intel Realsense D455 camera [22]. Data collection was performed in an indoor workspace with varied lighting conditions. The illumination levels and object distance were intentionally adjusted to capture emitter patterns with different sharpness and intensity, creating a varied dataset.
The initial dataset contained 10,000 paired images. To enhance model robustness and prevent overfitting, a data augmentation process was applied. For each image pair, five augmented versions were generated using random transformations, which included:
  • Rotation: Up to 30 degrees in either direction.
  • Flipping: Random horizontal and vertical flips.
  • Affine Transformations: Minor rotations and translations.
  • Zoom: Random zooming in by up to 30%.
  • Brightness and Contrast: Adjustments to simulate lighting variations.
The final data had 50000 paired sets of IR and greyscale RGB images and will be released as an open-sourced dataset.
III-D Composite Loss Function and Training
To produce high-quality image reconstructions, a composite loss function was used to guide the network’s training. This function is a weighted sum of multiple components, each designed to optimise a different aspect of the output, ensuring that the reconstructed image is not pixel accurate but perceptually and structurally similar to the ground truth. The main reason for choosing these losses besides mean absolute or mean squared error was due to the epipolar transformation of pixels between the IR and RGB images.
III-D1 Loss Components
The total loss is a weighted sum of several components. Let $\alpha,\beta,\gamma,\delta,\epsilon$, and $\zeta$ be the empirically tuned weighting coefficients for each corresponding loss component. The total loss is defined as:
[…]
  • Mean Absolute Error (MAE): This is a pixel-wise loss that minimises the average absolute difference between the predicted image $\hat{I}$ and the ground truth image $I$. It provides a foundational measure of pixel level accuracy.
    $\mathcal{L}_{mae}=\frac{1}{N}\sum_{i=1}^{N}|\hat{I}_{i}-I_{i}|$ (4)
  • Structural Similarity Index (SSIM) Loss: This loss measures the similarity in luminance, contrast, and structure between images. By minimising this loss, the network is encouraged to preserve the overall structural integrity of the scene [23].
    $\mathcal{L}_{ssim}=1-\text{SSIM}(\hat{I},I)$ (5)
  • High-Frequency and Sobel Edge Losses: These losses are used to preserve fine details and sharp boundaries. The high-frequency loss uses a Laplacian filter to penalise differences in the high-frequency components of the images, while the Sobel edge loss penalises differences in image gradients to ensure edges are accurately reproduced.
    $\mathcal{L}_{freq}=\left\|(\hat{I}-I)*K_{Laplacian}\right\|_{1}$ (6)
    $\mathcal{L}_{sobel}=\left\|\nabla_{x}\hat{I}-\nabla_{x}I\right\|_{1}+\left\|\nabla_{y}\hat{I}-\nabla_{y}I\right\|_{1}$ (7)
  • Perceptual Loss: This loss component uses a pre-trained VGG19 network as a feature extractor. The network’s weights, which were pre-trained on the ImageNet dataset [24], are frozen during training. This allows the model to leverage learned feature representations that are more aligned with human visual perception than simple pixel values. To apply this to greyscale images, the input is converted to a 3 channel format by concatenating the single channel three times. Instead of comparing pixels directly, this loss compares the feature maps from a set of specific intermediate layers within the VGG19 network. By using layers such as block3_conv2, block4_conv2, and the layers within block5, the loss encourages the reconstructed image to have similar high level content and style as the ground truth. This is crucial for generating an output that is visually convincing and suitable for high level vision tasks [25].
    $\mathcal{L}_{perceptual}=\sum_{j}\|\Phi_{j}(\hat{I})-\Phi_{j}(I)\|_{1}$ (8)
    where $\Phi_{j}(\cdot)$ is the feature map from layer $j$ of the VGG19 network.
  • Total Variation (TV) Loss: This regularizer promotes spatial smoothness and helps to reduce spurious noise in the output image [26].
    $\mathcal{L}_{tv}=\sum_{i,j}\left(\left(\hat{I}_{i+1,j}-\hat{I}_{i,j}\right)^{2}+\left(\hat{I}_{i,j+1}-\hat{I}_{i,j}\right)^{2}\right)$ (9)
III-D2 Weighting Strategy
The weighting coefficients for each component ($\alpha,\beta,\gamma,\delta,\epsilon,\zeta$) were experimentally tuned to balance their influence on the final output. The primary goal was to prioritise visual quality and structural integrity over a strict pixel-by-pixel match. High weights were assigned to $\mathcal{L}_{ssim}$ to preserve scene structure, while a significant but not overwhelming weight was given to $\mathcal{L}_{perceptual}$ to ensure the output was visually convincing. Moderate weights were used for $\mathcal{L}_{freq}$ and $\mathcal{L}_{sobel}$ to preserve fine details and edges while also not prioritising emitter patterns which can be classed as high-frequency patterns in the image. In contrast, very low weights were given to $\mathcal{L}_{mae}$ and $\mathcal{L}_{tv}$, as their purpose was to act as subtle regularizers rather than primary training objectives.
III-E Preprocessing and Training
After augmentation, the dataset, which contained 50,000 paired images, 90% of the data was split into a training set and the remaining 10% into the validation set. All images were resized to a consistent resolution of 640x480 pixels and normalised to the range of [0, 1].
The model was trained using the Adam optimizer with a CosineDecayRestarts learning rate schedule to minimise the composite loss function. To prevent overfitting, an early stopping mechanism with a patience of 30 epochs was implemented.
The model was trained for a total of 126 epochs, at which point it met the early stopping criteria, indicating that a sufficient level of convergence had been reached.
IV Experimental Results
This section presents a comprehensive evaluation of the proposed CLEAR-IR framework, focusing on its effectiveness in preparing infrared images for high-level computer vision tasks. Since this is the first work to specifically address the removal of active emitter patterns, the primary evaluation is a qualitative and quantitative comparison of the performance of enhancement techniques on the RGB stream versus the denoised images of the IR stream produced by this work.
IV-A Qualitative and Quantitative Analysis
To provide a visual assessment of the method’s reconstruction quality and its robustness in low light conditions, a qualitative comparison is first presented of a raw IR input, the model’s reconstructed output, and the corresponding ground truth image, as seen in Figure 4. The method’s output is also compared against state-of-the-art low-light RGB image enhancement techniques to highlight the inherent advantages of using the IR spectrum in these conditions.
(a) Raw IR Input
(b) Denoised Output
(c) Ground-Truth
Figure 4: A qualitative comparison of the raw IR input with emitter patterns, the reconstructed image, and the corresponding ground-truth image.
As seen in the visual examples, the reconstructed output successfully removes the emitter patterns while preserving key features such as object boundaries and textures. The output image is visually coherent and appears cleaner, making it a much more suitable input for downstream computer vision algorithms. Even object shadows which are hard to see in the raw IR image can be seen clearly in the reconstructed image.
The advantages of the CLEAR-IR method are further emphasised through a quantitative comparison with traditional methods such as Contrast Limited Adaptive Histogram Equalisation (CLAHE) [27], Retinex Theory inspired techniques [28] and state-of-the-art low-light RGB image enhancement techniques such as Zero-Reference Deep Curve Estimation (Zero-DCE) [29], Color and Intensity Decoupling Network (CIDNet) [30], and LLFormer [31]. For this comparison, the similarity of each method’s output to a ground truth RGB image is evaluated using the Structural Similarity Index Measure (SSIM).
The visual results in Figure 5 shows the outputs from different techniques on the scene. While other methods attempt to enhance the image, the enhanced RGB images are affected by the effects of noise observed in low-light scenes. The IR cameras are more robust to the effect of this noise in low-light, as a result, the CLEAR-IR framework produces a clean, artifact-free image that maintains the scene’s structural integrity, as reflected by its superior SSIM score. This demonstrates that the proposed method is more effective at generating a representation of the scene suitable for downstream vision tasks.
TABLE I: Quantitative comparison of image enhancement methods against ground truth.
Method SSIM
Retinex Theory [27] 0.7418
CLAHE [28] 0.5672
Zero-DCE [29] 0.7624
LLFormer [30] 0.7063
CIDNet [31] 0.7788
CLEAR-IR (Proposed Method) 0.9521
The results, presented in Table I, show that the denoised IR output achieves a significantly higher SSIM score and is capable of maintaining the true structural integrity of the scene, even when compared to methods that produce a colour output.
Beyond the general quality metrics, the true utility of the proposed CLEAR-IR framework is demonstrated by its ability to facilitate high-level robotics tasks that are otherwise compromised by the structured light patterns and noise inherent in raw IR images and RGB images respectively. Table II presents a comprehensive quantitative comparison of the performance of various image enhancement methods on three high level tasks such as ORB feature matching, object detection, and ArUco marker detection. The results highlight a critical distinction between methods that simply improve image aesthetics and those that produce a robust, clean representation suitable for downstream algorithms.
(a) Ground Truth RGB Scene
(b) Dark RGB Scene
(c) CLAHE
(d) Retinex Theory
(e) Zero-DCE
(f) LLFormer
(g) CIDNet
(h) CLEAR-IR (Proposed Method)
Figure 5: Visual comparison of a ground truth RGB scene, a dark RGB scene, and the outputs of various image enhancement methods, including the proposed Denoised IR method.
TABLE II: Performance comparison of image enhancement methods for various computer vision tasks.
Method ORB Matches (Lowe’s Test) Object Detection (Visible Objects) ArUco Marker Detection
Correct Incorrect Detected Missed Mislabelled Correct Missed
Retinex Theory [27] 49 451 1 9 0 5 1
CLAHE [28] 36 464 1 9 0 3 3
Zero-DCE [29] 61 439 0 10 0 5 1
LLFormer [30] 53 447 0 10 0 5 1
CIDNet [31] 66 434 1 9 0 6 0
Raw IR 31 469 0 9 1 0 6
CLEAR-IR (Proposed Method) 75 342 7 2 1 6 0
IV-B Feature Matching and Keypoint Detection
To assess the impact of emitter pattern removal on traditional computer vision algorithms, the performance of the Oriented FAST and Rotated BRIEF (ORB) [32] feature detector was evaluated. ORB is a robust algorithm for finding keypoints and feature descriptors, which are fundamental for tasks such as Simultaneous Localisation and Mapping (SLAM) and visual odometry.
As shown in Figure 6, the raw IR images exhibit numerous false keypoint detections due to the high-frequency nature of the emitter pattern. These spurious features lead to inaccurate and unreliable feature matches, which would compromise downstream tasks. In contrast, the denoised images from the CLEAR-IR model provide a clean representation where ORB is able to detect and match meaningful keypoints on actual scene structures, such as object boundaries and corners. This demonstrates that the proposed method effectively suppresses noise that could be misinterpreted as texture by feature detectors, thereby producing a more reliable input for foundational vision algorithms.
(a) ORB matches on raw IR image
(b) ORB matches on denoised IR image
Figure 6: ORB feature matching on (a) raw IR and (b) denoised IR images.
IV-C Object Detection and Segmentation
To demonstrate the utility of the denoised images for advanced perception tasks, a pre-trained YOLOv8 object detection and segmentation model was applied. This model was not trained on IR data, making this a zero-shot evaluation to determine if the proposed preprocessing step can make the images usable with standard RGB-based vision pipelines.
(a) YOLOv8 on RGB image
(b) YOLOv8 on raw IR image
(c) YOLOv8 on Retinex image
(d) YOLOv8 on CLEAR-IR image
Figure 7: Object detection and segmentation results of YOLOv8 on raw and denoised IR images.
On the raw IR images, YOLOv8 failed to detect or segment objects. The emitter patterns and resulting image noise prevented the model from recognising the distinct shapes and features of common objects. After preprocessing the images with the CLEAR-IR model, the performance of YOLOv8 dramatically improved. As illustrated in Figure 7, the model successfully detected and segmented objects in the denoised images, which were completely missed in the raw inputs. While tested on a variety of images, some labels were incorrect since the YOLOv8 model had not been trained on the specific objects present in the scene. For example, traffic cones were occasionally misclassified as bottles, and a white storage cabinet was misclassified as a refrigerator. Similar misclassifications were also observed in the RGB stream, indicating that this is primarily a limitation of the YOLOv8 model rather than the proposed CLEAR-IR framework. Importantly, despite label mismatches, the model was able to consistently detect and track these objects across frames, demonstrating the usability of CLEAR-IR outputs for object-level perception.
IV-D ArUco Marker Detection
ArUco markers are fiducial markers widely used in robotics for localisation and pose estimation. The ability to reliably detect these markers is crucial for robust navigation. However, structured light patterns can interfere with the marker’s distinct black-and-white grid, often leading to detection failures. This interference is clearly visible in the raw infrared (IR) image shown in Figure 8b. For comparison, a baseline for successful detection is provided by the standard RGB image, as seen in Figure 8a.
(a) Markers in the RGB image
(b) No markers in the raw IR image due to structured light interference
(c) Markers in the CIDNet output
(d) Markers in the CLEAR-IR output
Figure 8: Comparison of ArUco marker detection performance on RGB, raw IR, and denoised IR images.
The CLEAR-IR framework addresses these challenges by enabling robust ArUco marker detection in IR images. By removing the spurious dot patterns, the model ensures the marker’s grid is clearly visible and interpretable by the detection algorithm, as demonstrated in Figure 8d. This is a significant advancement for low-light scenes where raw IR images are typically unusable for this task. The denoised output allows for consistent and accurate marker identification, thus enabling reliable localisation and mapping. This outcome directly contrasts with findings in prior work, such as [33], which concluded that raw IR images were unsuitable for low-light tag detection without the use of retroreflective materials. Similarly, this approach simplifies methods like [34], which propose using specialised tags for low-light localisation, by requiring no special tags or materials.
V Conclusion
This work demonstrates that the CLEAR-IR framework effectively addresses a fundamental challenge in robotic perception by successfully removing structured light emitter patterns from infrared images. Leveraging a U-Net-based architecture, the method produces visually coherent and artifact free outputs that are compatible with existing computer vision pipelines in a zero-shot scenario. The significance of this contribution is further evidenced by the substantial improvement in performance for crucial downstream tasks, including feature matching, object detection, and marker detection. As confirmed by quantitative results, this approach establishes a new standard, demonstrably surpassing state-of-the-art low-light RGB enhancement techniques. The CLEAR-IR framework thus provides a robust and reliable foundation for enabling autonomous systems to operate effectively in a wider range of challenging lighting conditions.
VI Future Direction
Future work will concentrate on two primary objectives to advance the CLEAR-IR framework’s utility in practical applications such as enhancing computational efficiency and enabling robust integration with visual Simultaneous Localisation and Mapping (vSLAM) systems. To improve processing time, research will explore model optimisation techniques, such as network quantisation and model pruning.
For integration with vSLAM, future work will focus on assessing the effectiveness of CLEAR-IR within a complete visual SLAM pipeline. This will involve real-time deployment and a quantitative evaluation of its impact on trajectory estimation accuracy and loop closure detection in illumination-constrained environments.
Acknowledgment
RAICo contributed to this work by providing funding and a supportive research environment. Jongyun Kim assisted the main author with setup and data collection in a dark environment.